Monica is one year older than twice her sister's age. The difference of their ages is 18. Find theor present ages.
Answer by mananth(16946)

monica x years
sister y years
...
x=2y+1
x-2y=1................1
x-y = 18..............2
multiply equation by -1 and add it up to equation 2
-x+2y=-1
x-y = 18
y=17 sister's age
...
x-y 1=8
x=18+y
x=18+17 = 35 monica'a age
......
OR
if sister is x years
monica = 2x+1
...
2x+1-x=18
x= 18-1
x= 17 years sister's age
2x+1= 2*17+1= 35 monica's age